Scilab 1.0 On line doc

Scilab Group

February 13, 1998

Chapter 1

Test

1.0.1 errcatch error trapping

CALLING SEQUENCE :

errcatch(n [,'action'] [,'option'])

PARAMETERS :

n : integer

action, option : strings

DESCRIPTION :

errcatch gives an ”action” (error-handler) to be performed when an error of type n occurs. n has the followin meaning:

if n>0, n is the error number to trap if n<0 all errors are to be trapped

action is one of the following character strings:

"pause" : a pause is executed when trapping the error. This option is useful for debugging purposes. "continue" : next instruction in the function or exec files is executed, current instruction is ignored. This

option is useful for error recovery. "kill" : default mode, all intermediate functions are killed, scilab goes back to the level 0 prompt. "stop" : interrupts the current Scilab session (useful when Scilab is called from an external program).

option is the character string 'nomessage' for killing error message.

See Also : errclear ??, iserror ??

Index^{E_{errcatch, 1}}

2